OpenStudy (anonymous):

the legend of a map says that .5 inches is equal to 10 miles on a map. if a highway on a map is 1.75 inches long, how long is the actual highway

OpenStudy (anonymous):

You are comparing inches to miles... so set up a proportion: inches is to miles...as inches is to miles so your equation would be .5/10 = 1.75/x (x is what you are looking for).

OpenStudy (mertsj):

I see the asker has gone offline.

OpenStudy (anonymous):

yep, easyaspi is correct, however you might wanna set up the ratios in reverse to make it easier to solve for x... 10miles/.5 inches = x miles/1.75 inches... so x=(1.75)(10/.5) miles :o)